Cash flow forecasts - case study - cash flow management

Wholesale Phones Ltd

Three shareholders, Sandra, Sumira and Benson, have just set up a small company that buys mobile phones from Asia and the sells them to wholesalers in Europe. Marketing and selling is done using the Internet. When they started operations, Benson the Accountant, arranged an overdraft line with the Bank for $25,000. They had also recently deposited $20,000 in cash in return for the initial share issue (equity).

The early weeks were very hard, but the market was growing. Sales were very good, and all the owners' efforts were put into buying and selling. Credit was ignored, and Benson let the work on the cash flow forecast slip. The owners are in the boardroom when the receptionist interrupts them.

'Benson, phone the Bank Manager at once. Unless you talk to her, she will call our overdraft in.'

'What is she on about', said Sandra, 'does that mean I will have to cancel the order for my BMW?'

Benson went to his office. He had not really checked the cash position for two months.

He was a little worried. His records showed the following.

Opening bank balance - $2,500 credit

Overhead expenses

Rent: Office $5,000 per month. Two months paid.
Furniture: $1,000 per month. One month paid.
Cars: $2,500 per month. One month paid.
Insurance: $2,000 per month. Two months paid.
Tax: $500 per month. None paid.
Services: $250 per month. Two months paid.
Salaries and benefits: $12,000 per month. Two months paid.
Misc.: $1,000 per month. One month paid.

Sales                                               Purchases

Period 1: $15,000 All paid for.                           Period 1: $10,000 Paid.                         
Period 2: $24,000 All paid for.                           Period 2: $18,000 Paid.
Period 3: $40,000 Half paid for.                         Period 3: $25,000 Paid.


Was he right to be worried, and will Sandra get her car? He had set up an overdraft limit of $25,000 with the bank. How big a line might be needed?
Answer #1

Amounts are in $

Opening bank balance = 2,500

Expenses during the two months

Rent = 5,000 x 2 = 10,000

Furniture = 1,000 (another 1,000 payable)

Car = 2,500 (another 2,500 payable)

Insurance = 2,000 x 2 = 4,000

Tax = 0 (1,000 payable)

Services = 250 x 2 = 500

Salaries = 12,000 x 2 = 24,000

Miscellaneous = 1,000 (another 1,000 payable)

Total expenses paid = 43,000

Total payable expenses = 5,500

Receipts from Sales = 15,000 + 24,000 + 20,000 = 59,000

(Another 20,000 is yet to be received)

Payment for purchases = 10,000 + 18,000 + 25,000 = 53,000

Present Overdraft balance = Opening balance - Expenses paid - Purchase payments + Sales Receipts

= 2,500 - 43,000 - 53,000 + 59,000

= -34,500

So overdraft is having a balance of 34,500 which is in exces of the authorised limit. This may result in bank charging higher interest and calling in overdraft by the bank.

(Assumed as the initial cash deposit of $20,000 through share issue is already considered in getting the opening bank balance of $2,500. If we assume otherwise, the Overdraft balance would have been $14,500 debit and there are very less chances for the bank to call in Overdraft, So we assumed the first option that means the initial deposit is already considered in getting that opening balance of $2,500)

It is right to be worried as there exists two threats, higher interest charging by bank and also the threat of call in of overdraft.

After getting the remaining 20,000 receipts from Sales (which are only half paid as of now) it will be first used to settle the payable expenses (which amount to 5,500) and this will lead to revision of Overdraft balance to 20,000 debit (-34,500 - 5,500 + 20,000). This will bring down the overdraft balance below authorised limit and helps in avoiding higher rate of interest. But even now this balance will not be sufficient to buy a BMW. For this to happen we have to increase our line of credit with the bank, this depends on the Car Price (Cash Price or down payment and Installment amounts) or he has to get a separate loan for this purpose.